Circular No. 7692

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

NOVA SAGITTARII 2001 No. 2 Alfredo J. S. Pereira, Cabo da Roca, Portugal, reports his visual dis-

covery of an apparent nova (mag 7.6) on Aug. 26.866 UT during a regular patrol with 14×100 binoculars; he did not notice anything new during scans of the same region to m_{v} ∼ 7.5–8 on Aug. 25.95 and to m_{v} ∼ 8.5 on Aug. 21.9 and 22.96. Following a request by the Central Bureau, C. Jacques, Belo Horizonte, Brazil, sends the following precise position for the new star from CCD frames taken by G. Nappi on Aug. 27.00 (mag 9.5): α = 18^{h}24^{m}46^{s}.04, δ = −30^{o}00′41′′.1 (equinox 2000.0). A. Hale, Cloudcroft, NM, reports CCD position end figures 46^{s}.06, 40′′.9, adding that nothing is visible at this location on the Digitized Sky Survey image of this field. P. M. Kilmartin, Mt. John Observatory, University of Canterbury, reports the following photoelectric photometry obtained with the 0.6-m f/16 reflector (reference stars Cousins E745 and E746; uncertainties < 0.01 mag in all measures; airmass 1.03–1.04): Aug. 27.339, V = 7.46, U − B = −0.71, B − V = +0.55, V − R = +0.53, V − I = +0.99; 27.353, 7.51, −0.74, +0.56, +0.55, +1.01; 27.369, 7.58, −0.75, +0.55, +0.56, +1.04; 27.375, 7.59, −0.76, +0.56, +0.56, +1.04. Further visual magnitude estimates (those by Scovil and Griese sent via J. Mattei, AAVSO): Aug. 26.913, 7.3 (Pereira); 26.930, 7.4 (Pereira); 26.940, 7.4 (Pereira); 26.967, 7.2 (Pereira); 26.996, 7.2: (Pereira); 27.10, 6.4 (C. Scovil, Stamford, CT); 27.10, 6.6 (J. Griese, Rocky Hill, CT); 27.18, 7.4 (Hale); 27.38, 8.0 (A. Jones, Nelson, New Zealand). On Aug. 27.10, M. Mattei (Littleton, MA) reports V = 7.29 on a CCD exposure taken with a 135-mm camera lens.

SUPERNOVA 2001dx IN PGC 63222 W. D. Li, University of California at Berkeley, reports the discovery

by LOTOSS (cf. IAUC 7514) of an apparent supernova (mag ∼ 17.2) in unfiltered images taken with the Katzman Automatic Imaging Telescope (KAIT) on Aug. 26.3 and 27.3 UT. The new object is located at α = 19^{h}27^{m}16^{s}.83, δ = +65^{o}18′58′′.8 (equinox 2000.0), which is 5′′.4 west of the nucleus of PGC 63222. A KAIT image taken on Aug. 17.3 showed nothing at this position (limiting mag ∼ 19.5).

SUPERNOVA 2001du IN NGC 1365 CCD position by Jacques for SN 2001du (cf. IAUC 7690): Aug. 26.22

UT, mag 15.5, α = 3^{h}33^{m}29^{s}.11, δ = −36^{o}08′32′′.5 (equinox 2000.0).

2001 August 27  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green